# When Representations Align: Universality in Representation Learning Dynamics

### v1

@@ 1 Introduction
Studies have shown that although the exact representation between different neural networks may differ, some important aspects of the computational structure they learn are universal (Maheswaranathan et al., 2019; McMahan et al., 2021; Huang et al., 2021). Even for certain areas in the brain, similarities in representational structure have been found with deep neural networks trained on natural image data (Yamins et al., 2014; Yamins & DiCarlo, 2016; Saxe et al., 2021; Conwell et al., 2023; Bracci et al., 2023). The data used in learning may play a larger role than the implementation details of the model, a fact which has been empirically demonstrated in vision transformers and convolutional neural networks (Bouchacourt et al., 2021; Conwell et al., 2023) and studied theoretically (Goldt et al., 2020). Additionally, neural scaling laws describe a formal relationship between model or dataset size and performance that appears to hold without making direct reference to architecture (Kaplan et al., 2020; Hoffmann et al., 2022). These observations together license universal theories: mathematical descriptions of learning that sidestep certain implementation details in order to model behavior common among architectures.
@@
Here we propose an approach to modelling universal representation learning, aiming to provide conceptual explanations for phenomena universally observed in learning systems. Our main contributions are as follows:
@@ 1.1 Related Work
A notable exception in terms of universal solutions is that of the neural tangent kernel literature (Lee et al., 2019; Jacot et al., 2020; Golikov et al., 2022), which does provide exact solutions applicable to a wide range of models. The limitation here is that the neural tangent kernel solutions are only valid for infinitely wide neural networks at a specific weight initialization scale (Woodworth et al., 2020). In particular, the initialization regime used is one of relatively high weights, where the dataset does not impact the representation (Chizat et al., 2020; Flesch et al., 2022). We seek a similar universal theory, applicable to the feature learning regime, where the dataset controls the representation.
@@
A large body of work has investigated gradient descent as a source of generalization performance in deep neural networks, as opposed to explicit regularization and inductive biases in the architecture (Neyshabur et al., 2015; Gunasekar et al., 2018; Chizat & Bach, 2020; Soudry et al., 2022). This work is mainly either empirical or focused on analyses for examples of specific architectures. Here instead we consider a modeling approach, aiming to provide intuition into the universal nature of implicit bias and observed similarities between representational structures across architectures.
@@ 2.1 Function Dynamics
Since we are interested in studying the dynamics of the representational space, we will split $f_{\theta}$ into two smooth maps, an encoder map $h_{\theta}:X\to H$ which assigns to each input in $X$ a representation in some intermediate layer $H$ in the neural network, and a decoder map $y_{\theta}:H\to Y$ which assigns a predicted output to each possible representation in $H$. We may then write the loss in Equation 1 as:
$L(h_{\theta},y_{\theta})=\frac{1}{2}\langle||y_{\theta}(h_{\theta}(x_{i}))-y_{i}||^{2}\rangle_{\mathcal{D}}.$ (2)
The dynamics on these maps follow from applying the gradient descent update rule to the parameters $\theta$ defining them:
$\Delta\theta=-\eta\nabla_{\theta}L,$ (3)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\theta=-\frac{1}{\tau}\nabla_{\theta}L,$ (4)
where $\tau:=\frac{1}{\eta}$. The problem with Equation 4 is that it explicitly depends on the parametrization of the neural network, whereas a universal approach requires ignoring such implementation details. Ideally we would like to optimize the loss $L$ directly with respect to the maps $h_{\theta}$ and $y_{\theta}$, since if the model has high expressivity, i.e. the architecture is large and complex enough, perhaps it has enough freedom to behave as an arbitrary smooth map. However, it is unclear how this can be achieved mathematically [^1] One may be tempted to try replacing the right hand side of Equation 4 with functional derivatives of $L$ with respect to $h_{\theta}$ and $y_{\theta}$, however this will result in a term involving the gradient of $y_{\theta}$, and the dynamics of this gradient explicitly depend on $\theta$..
@@ 2.2 Two Point Interaction
In general, the modelling issue of parametrization dependent dynamics does not appear to have a resolution. So let us instead restrict ourselves to a simpler case, where the only property we do know about the unspecified network, the smoothness of the maps $h_{\theta}$ and $y_{\theta}$, can be exploited. During training the representations of different datapoints move around in the hidden space $H$, and may get near each other and interact. Understanding such an interaction on its own may already provide some insight into the representational learning behavior of the model on the full dataset.
To model this two point interaction, let us consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2})\}$. If the representations $h_{\theta}(x_{1})$ and $h_{\theta}(x_{2})$ are close enough, then by smoothness it is reasonable to a take a linear approximation of $h_{\theta}$ and $y_{\theta}$ around the mean of the two points:
$\begin{split}h_{\theta}(x_{i})&=h_{\theta}(\frac{x_{2}+x_{1}}{2})+\frac{1}{2}D_{h}(x_{i}-x_{\neg i})\\ y_{\theta}(x_{i})&=y_{\theta}(h_{\theta}(\frac{x_{2}+x_{1}}{2}))+\frac{1}{2}D_{y}(D_{h}(x_{i}-x_{\neg i}))\end{split},$ (5)
where $D_{h}$ and $D_{y}$ are the Jacobian matrices of $h_{\theta}$ and $y_{\theta}$ respectively.
Presumably, as gradient descent optimizes $\theta$ in order to reduce the loss, at the two point interaction this will effectively optimize the linearization parameters $D_{h}$, $D_{y}$ and $\bar{y}:=y_{\theta}(h_{\theta}(\frac{x_{2}+x_{1}}{2}))$ in order to minimize the loss [^2] Potentially $h_{\theta}(\frac{x_{2}+x_{1}}{2})$ should also be considered as a linearization parameter, but it will turn out to not play a role in the following analysis.. Assuming expressivity, i.e. “enough freedom” in the architecture such that there are no additional constraints on the dynamics, we can model the dynamics by applying gradient descent directly to the linearization parameters, i.e.
@@
Note that these dynamics are not necessarily equal to the actual dynamics on the linearization parameters induced by the model parameters $\theta$. Equation 6 is the main modelling assumption, and intended as an effective theory in the regime of large complex architectures, where behavior is unconstrained by their parametrization. The details of how $h_{\theta}$ and $y_{\theta}$ are parameterized determine the effective learning rates $1/\tau_{\bar{y}},1/\tau_{h}$ and $1/\tau_{y}$, but are otherwise not present in the effective theory.
@@ 3.1 Comparison of the Dynamics
In Figure 2 we compare the learning dynamics of a neural network (20 layer, leaky ReLU, 500 units per layer) trained on a two point dataset to numerical solutions of the effective theory, as well as variations of this network with different architectures and nonlinearities. Despite only needing to fit two constants, the theory matches all three curves quite well for all networks.
@@ 3.2 Varying Initial Weights
For small initial weights, the variables $||dh||,||dy||$ and $w$ start small, so the right hand side of Equation 7 must be small too. Therefore, the dynamics are initially stagnant resulting in plateau-like behavior, whereas for large initial weights we have immediate exponential decay, as derived analytically in Section A.5. Such a transition from an initial plateau to immediate decay is observed experimentally, as can be seen in Figure 3 (top). Going to even larger initial weights, the representations start off far apart, so the linear approximation breaks down and the effective theory no longer matches the network.
@@ 3.4 Larger Datasets
So far, we have only considered the interaction of two points in isolation. Real datasets consist of many points, which may influence the behavior two interacting datapoints exhibit. To investigate the validity of the theory in this setting, we trained a model on the MNIST dataset, and tracked two datapoints during learning. As can be seen in Figure 4, the resulting dynamics still resemble the theory reasonably well.
Figure 4: Learning dynamics of two tracked points from different classes in MNIST compared against the theory after fitting two constants. The deep neural network has 4 fully connected layers, 100 units per layer, leaky ReLU activation and is initialized at small weights.
@@ 4.1 Lazy and Rich Learning Regimes
This separation between a random regime and a structured regime closely resembles the phenomenon of the “rich” and “lazy” learning that deep neural networks exhibit (Chizat et al., 2020; Flesch et al., 2021, 2022; Atanasov et al., 2022). In particular, a crucial factor determining the regime is the initial weight scale (Woodworth et al., 2020).
@@ 4.2 Intuition
There is an intuition as to why smoothness constraints can result in these two different representational regimes. At large weights, representations are initially far apart, so the flexibility of the network allows the decoder to individually learn the correct output freely for each datapoint, without need of significantly adjusting the representational structure. The dominating structure in the final representation is determined by the structure already present at initialization.
At small weights, representations are nearby and thus, due to smoothness constraints, they can no longer freely learn outputs without influencing each other. To illustrate this, consider the case where we have two datapoints with the same target output but only one already has the correct output prediction. There are two ways to learn the correct output for the other datapoint: (1) the decoder map can adjust at its representation and (2) the encoder map can move the representation towards the representation of the other datapoint, which, since that already has the correct output, will move the output of the first datapoint to the correct output, as the decoder map is continuous. The second option therefore causes an effect on representations, where they move closer when their outputs agree. The opposite effect also holds: when two datapoints have different target outputs but the same representation, by continuity of the decoder map, they cannot both have the correct output prediction. Their representations must move apart to fit the data.
@@
4.3 Rich Structure
The representation at small initial weights is of particular interest, as it appears to be based on a structure found in the dataset, rather than formed by the random weight initialization. Note that again the only unknown quantities here are the two effective learning rates. By fitting these to the learning trajectories we can make predictions about the final representational, as shown in Figure 5. Although the prediction here is only for a single pair of datapoints, given a simple enough dataset, we can attempt to extrapolate the result to larger datasets.
@@ 4.6 MNIST
When dealing with more complex datasets we cannot always apply Equation 10 directly. This is because it only predicts an interaction between two datapoints. Even simple geometric effects such constraints coming from the triangle inequality are not respected. Nevertheless, in Figure 9 we see that part of the representational structure still resembles the theory when a network is trained on the MNIST dataset. Here we applied a simple method to reduce violations of the triangle inequality, as detailed in Section B.4.
Figure 9: Representational structure on the MNIST dataset at small initial weights. Pairwise distances of the first 100 datapoints in MNIST after averaging over 50 trials of a 4-layer, leaky-ReLU, fully connected network trained on the full dataset (left) compared to the theory after exponential weighing (right). The Pearson correlation coefficient between the theory and experiment is 0.703.
@@ 6 Conclusion
We have introduced an effective theory for the universal part of the learning dynamics common amongst different neural network architectures, and shown it already exhibits structured representational learning, in particular at small initial weights. The aim here is not so much to point to this specific theory as an accurate universal model, rather it is to suggest that some of the essentials required for representation learning may already be contained within gradient descent, as opposed to solely coming from inductive biases contained in the choice of architecture. Therefore, perhaps a more universal perspective to learning dynamics is warranted. Additionally, the effective theory underlines the initial weight scale as a crucial factor for the formation of the final representational structure.
A significant limitation to this specific approach for modelling universality is that there is no clear path to extending it to larger datasets. More work is needed to find a satisfying resolution to dealing with more complex data. Additionally, many architectures do have inductive biases that affect the learned representation, and the theory here does not take those into account. The representational effect modelled here may in fact interact with these inductive biases, causing potentially interesting effects on the representation that are worth exploring.
@@ B.2 XOR Task
MINST leaky relu 0.5 0.002 500 100 4
@@ B.4 Experiment on MNIST
The full training set of MNIST handwritten digit classification was used during training. The representational structure was plotted for the first 100 digits, sorted by digit. For the learning dynamics, the first occurrence of the zero digit in MNIST and the first one digit were tracked.
@@
The architecture used is a fully connected, leaky ReLU network, with only 4 hidden layers instead of the usual 20 to allow for faster convergence in the feature learning regime. The same hyperparameters were used for both the learning dynamics and representational structure experiments and can be found in Table 2 (bottom).
@@
A prediction for pairwise representational distances does not necessarily directly translate into a prediction for the full dataset representation. Even when assuming pairwise distances in the full dataset behave independently, the two datapoint prediction may end up breaking the triangle inequality, i.e. it may predict that representations of datapoints A and B are close, as well as B and C, but still predict A and C to be far away.

### v2

@@ 1 Introduction
Studies have shown that although the exact representation between different neural networks may differ, some important aspects of the computational structure they learn are universal (Maheswaranathan et al., 2019; McMahan et al., 2021; Huang et al., 2021). Even for certain areas in the brain, similarities in representational structure have been found with deep neural networks trained on natural image data (Yamins et al., 2014; Yamins & DiCarlo, 2016; Saxe et al., 2021; Conwell et al., 2023; Bracci et al., 2023). The data used in learning may play a larger role than the implementation details of the model, a fact which has been empirically demonstrated in vision transformers and convolutional neural networks (Bouchacourt et al., 2021; Conwell et al., 2023) and studied theoretically (Goldt et al., 2019). Additionally, neural scaling laws describe a formal relationship between model or dataset size and performance that appears to hold without making direct reference to architecture (Kaplan et al., 2020; Hoffmann et al., 2022). More recently, it has been found that an alternative architecture performs roughly the same as the transformer, given a similar parameter count (Wang et al., 2022). These observations together license universal theories: mathematical descriptions of learning that sidestep certain implementation details in order to model behavior common among architectures.
@@
Here we propose an approach to modelling universal representation learning, striving to provide conceptual explanations for phenomena universally observed in learning systems. We seek to summarize known results in deep learning within a single framework. Our main contributions are as follows:
@@ 1.1 Related Work
A notable exception in terms of universal solutions is that of the neural tangent kernel literature (Lee et al., 2019; Jacot et al., 2018; Golikov et al., 2022), which does provide exact solutions applicable to a wide range of models. The limitation here is that the neural tangent kernel solutions are only valid for infinitely wide neural networks at a specific weight initialization scale (Woodworth et al., 2020). In particular, the initialization regime used is one of relatively high weights, where the dataset does not impact the representation (Chizat et al., 2019; Flesch et al., 2022). We seek a similar universal theory, applicable to the feature learning regime, where the dataset controls the representation.
@@
A large body of work has investigated gradient descent as a source of generalization performance in deep neural networks, as opposed to explicit regularization and inductive biases in the architecture (Neyshabur et al., 2015; Gunasekar et al., 2018; Chizat & Bach, 2020; Soudry et al., 2018). This work is mainly either empirical or focused on analyses for examples of specific architectures. Here instead we consider a modeling approach, aiming to provide intuition into the universal nature of implicit bias and observed similarities between representational structures across architectures.
@@
Physics inspired approaches.
The approach to modelling a two point interacting between datapoints taken here is inspired by intuitions from physics surrounding interacting particles. Physics-inspired particle approaches are prominent in the deep learning theory (Seung et al., 1992; Rotskoff & Vanden-Eijnden, 2018a, b; Baek et al., 2024), most notably the application of mean field theory (Saul et al., 1996; Poole et al., 2016; Xiao et al., 2018; Mei et al., 2018, 2019; Pandey & Xu, 2019).
@@ 2.1 Function Dynamics
Since we are interested in studying the dynamics of the representational space, we will split $f_{\theta}$ into two smooth maps, an encoder map $h_{\phi}:X\to H$ which assigns to each input in $X$ a representation in some intermediate layer $H$ in the neural network, and a decoder map $y_{\psi}:H\to Y$ which assigns a predicted output to each possible representation in $H$. We may then write the loss in Equation 1 as:
$L(h_{\phi},y_{\psi})=\frac{1}{2}\langle||y_{\psi}(h_{\phi}(x_{i}))-y_{i}||^{2}\rangle_{\mathcal{D}}.$ (2)
The dynamics on these maps follow from applying the gradient descent update rule to the parameters $\phi$ and $\psi$ defining them:
$\Delta\phi=-\eta\nabla_{\phi}L\quad\text{and}\quad\Delta\psi=-\eta\nabla_{\psi}L,$ (3)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\phi=-\frac{1}{\tau}\nabla_{\phi}L\quad\text{and}\quad\frac{\mathrm{d}}{\mathrm{d}t}\psi=-\frac{1}{\tau}\nabla_{\psi}L,$ (4)
where $\tau:=\frac{1}{\eta}$. The problem with Equation 4 is that it explicitly depends on the parametrization of the neural network, whereas a universal approach requires ignoring such implementation details. Ideally we would like to optimize the loss $L$ directly with respect to the maps $h_{\phi}$ and $y_{\psi}$, since if the model has high expressivity, i.e. the architecture is large and complex enough, perhaps it has enough freedom to behave as an arbitrary smooth map. However, it is unclear how this can be achieved mathematically [^1] One may be tempted to try replacing the right hand side of Equation 4 with functional derivatives of $L$ with respect to $h_{\phi}$ and $y_{\psi}$, however this will result in a term involving the gradient of $y_{\psi}$, and the dynamics of this gradient explicitly depend on $\psi$..
@@ 2.2 Two Point Interaction
In general, the modelling issue of parametrization dependent dynamics does not appear to have a resolution. So let us instead restrict ourselves to a simpler case, where the only property we do know about the unspecified network, the smoothness of the maps $h_{\phi}$ and $y_{\psi}$, can be exploited. During training the representations of different datapoints move around in the hidden space $H$, and may get near each other and interact. Understanding such an interaction on its own may already provide some insight into the representational learning behavior of the model on the full dataset.
To model this two point interaction, let us consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2})\}$. If the representations $h_{\phi}(x_{1})$ and $h_{\phi}(x_{2})$ are close enough, then by smoothness it is reasonable to a take a linear approximation of $h_{\phi}$ and $y_{\psi}$ around the mean of the two points:
$\begin{split}h_{\phi}(x_{i})&\approx h_{\phi}(\frac{x_{2}+x_{1}}{2})+\frac{1}{2}D_{h}(x_{i}-x_{\neg i})\\ y_{\psi}(x_{i})&\approx y_{\psi}(h_{\phi}(\frac{x_{2}+x_{1}}{2}))+\frac{1}{2}D_{y}(D_{h}(x_{i}-x_{\neg i}))\end{split},$ (5)
where $D_{h}$ and $D_{y}$ are the Jacobian matrices of $h_{\phi}$ and $y_{\psi}$ respectively.
Presumably, as gradient descent optimizes $\phi$ and $\psi$ in order to reduce the loss, at the two point interaction this will effectively optimize the linearization parameters $D_{h}$, $D_{y}$ and $\bar{y}:=y_{\psi}(h_{\phi}(\frac{x_{2}+x_{1}}{2}))$ in order to minimize the loss [^2] Potentially $h_{\phi}(\frac{x_{2}+x_{1}}{2})$ should also be considered as a linearization parameter, but it will turn out to not play a role in the following analysis.. Assuming expressivity, i.e. “enough freedom” in the architecture such that there are no additional constraints on the dynamics, we can model the dynamics by applying gradient descent directly to the linearization parameters, i.e.
@@
Note that these dynamics are not necessarily equal to the actual dynamics on the linearization parameters induced by the model parameters $\phi$ and $\psi$. Equation 6 is the main modelling assumption, and intended as an effective theory in the regime of large complex architectures, where behavior is unconstrained by their parametrization. The details of how $h_{\phi}$ and $y_{\psi}$ are parameterized determine the effective learning rates $1/\tau_{\bar{y}},1/\tau_{h}$ and $1/\tau_{y}$, but are otherwise not present in the effective theory.
In fact, if one were to interpret the linearization parameters as model weights, then the effective theory dynamics are mathematically the same as that of a two layer linear network with bias and differing learning rates for each of the layers. However, we are not using a two layer linear network to model the neural network here, instead we are locally modeling a two point interaction.
@@ 3.1 Comparison of the Dynamics
In Figure 2 we compare the learning dynamics of a neural network (20 layer, leaky ReLU, 500 units per layer) trained on a two point dataset to numerical solutions of the effective theory, as well as variations of this network with different architectures and nonlinearities. Despite only needing to fit two constants, the theory matches all three curves quite well for all networks. Note that this is not solely a result of having two degrees of freedom (see Section C.1).
@@ 3.2 Varying Initial Weights
For small initial weights, the variables $||dh||,||dy||$ and $w$ start small [^3] There is an assumption here about the architecture. In principle, small weights could mean anything if nothing is know about how the smooth maps are parameterized, however, for typical neural network architectures these variables are small. , so the right hand side of Equation 7 must be small too. Therefore, the dynamics are initially stagnant resulting in plateau-like behavior, whereas for large initial weights we have immediate exponential decay, as derived analytically in Section A.5. Such a transition from an initial plateau to immediate decay is observed experimentally, as can be seen in Figure 3 (top). Going to even larger initial weights, the representations start off far apart, so the linear approximation breaks down and the effective theory no longer matches the network.
@@ 3.4 Larger Datasets
We have only considered the interaction of two points in isolation. Real datasets consist of many points, which may influence the behavior two interacting datapoints exhibit. To investigate the validity of the theory in this setting, we trained a model on the MNIST dataset, and tracked two distinguishable datapoints. The resulting dynamics (Figure 4) still qualitatively resemble the theory.
Figure 4: Learning dynamics of a randomly selected zero digit and one digit in MNIST compared against the theory after fitting two constants. The deep neural network has 4 fully connected layers, 100 units per layer, leaky ReLU activation and is initialized at small weights.
@@ 4.1 Lazy and Rich Learning Regimes
This separation between a random regime and a structured regime closely resembles the phenomenon of the “rich” and “lazy” learning that deep neural networks exhibit (Chizat et al., 2019; Flesch et al., 2021, 2022; Atanasov et al., 2022). In particular, a crucial factor determining the regime is the initial weight scale (Woodworth et al., 2020).
4.2 Rich Structure
The representation at small initial weights is of particular interest, as it appears to be based on a structure found in the dataset, rather than formed by the random weight initialization. Note that again the only unknown quantities here are the two effective learning rates. By fitting these to the learning trajectories at a fixed choice of input and target output distances of the datapoints, we can extrapolate to predictions about the final representational distance for other distances, as shown in Figure 5.
@@ 4.3 Intuition
There is an intuition as to why smoothness constraints can result in these two different representational regimes. At large weights, representations are initially far apart, so the flexibility of the network allows the decoder to individually learn the correct output freely for each datapoint, without need of adjusting the representational structure. The dominating structure in the final representation is determined by the structure already present at initialization.
At small weights, representations are nearby and thus, due to smoothness constraints, they can no longer freely learn outputs without influencing each other. To illustrate this, consider the case where we have two datapoints with the same target output but only one already has the correct output prediction. There are two ways to learn the correct output for the other datapoint: (1) the decoder map can adjust at its representation and (2) the encoder map can move the representation towards the representation of the other datapoint, which due to continuity of the decoder map will align its output to that of the other datapoint, which already has the correct output. The second option therefore causes an effect on representations, where they move closer when their outputs agree. The opposite effect also holds: when two datapoints have different target outputs but the same representation, by continuity they cannot both have the correct output prediction. Their representations must move apart to fit the data.
@@ 4.6 MNIST
When dealing with more complex datasets we cannot always apply Equation 10 directly. This is because it only predicts an interaction between two datapoints. Even simple geometric constraints such as the triangle inequality are not expected. Nevertheless, in Figure 9 we see that part of the representational structure still resembles the theory when a network is trained on the MNIST dataset. Here we applied a simple method to reduce violations of the triangle inequality, as detailed in Section B.4.
Figure 9: Representational structure on the MNIST dataset at small initial weights. Pairwise distances of the first 100 datapoints in MNIST after averaging over 50 trials of a 4-layer, leaky-ReLU, fully connected network trained on the full dataset (left) compared to the theory after exponential weighing (right). The Pearson correlation coefficient between the theory and experiment is 0.703 and when not factoring in the diagonal blocks it is 0.512.
@@ 6 Conclusion
We have introduced an effective theory for the universal part of the learning dynamics common amongst different neural network architectures, and shown it already exhibits structured representational learning, in particular at small initial weights. The aim here is not so much to point to this specific theory as an accurate universal model, rather it is to suggest that some of the essentials required for representation learning may already be contained within gradient descent, as opposed to solely coming from inductive biases contained in the choice of architecture. Therefore, perhaps a more universal perspective to learning dynamics is possible. Additionally, the effective theory underlines the initial weight scale as a crucial factor for the formation of the final representational structure.
A significant limitation to this specific approach for modelling universality is that there is no clear path to extending it to larger datasets. More work is needed to find a satisfying resolution to dealing with more complex data. Additionally, many architectures do have inductive biases that affect the learned representation, and the theory here does not take those into account. The representational effect modelled here likely interacts with these inductive biases, causing potentially interesting effects on the representation that are worth exploring.
Impact Statement
This paper presents work whose goal is to increase understanding of deep learning, which may lead to advancements in the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.2 XOR Task
MINST Trajectories leaky relu 0.5 0.003 200 200 4
MINST Structure leaky relu 0.5 0.005 200 200 4
@@ B.4 Experiment on MNIST
The full training set of MNIST handwritten digit classification was used during training. The representational structure was plotted for the first 100 digits, sorted by digit. For the learning dynamics, a zero digit and a one digit were selected at random. The digits zero and one were chosen since they look fairly distinct, compared to e.g. three and eight. Trajectories look slightly worse for more similar digits, but still roughly share a similar shape to the effective theory.
@@
The architecture used is a fully connected, leaky ReLU network, with only 4 hidden layers instead of the usual 20 to allow for faster convergence in the feature learning regime. The hyperparameters used for both the learning dynamics and representational structure experiments can be found in Table 2 (bottom). Initial weight parameters were chosen to display rich learning behaviour, while still converging within reasonable time.
Representational Structure
For comparing the final representational structure to the theory prediction, the model was trained 50 times at varying random seeds. The resulting representational distances between the first 100 digits in the dataset were averaged. The Pearson correlation coefficient is computed between the vector of all pairwise final representational distances and their respective theory predictions from Equation 10.
@@
Applying Equation 10 directly makes little sense, as a prediction for pairwise representational distances does not necessarily directly translate into a prediction for the full dataset representation. Even when assuming pairwise distances in the full dataset behave independently, the two datapoint prediction may end up breaking the triangle inequality, i.e. it may predict that representations of datapoints A and B are close, as well as B and C, but still predict A and C to be far away. Thus, we must apply a preprocessing step to take the geometry into account.